OpenStudy (ittybitty14):

FAN + MEDAL + TESTIMONIAL Annie is a financially responsible person who keeps a detailed record of the activities in her bank account. (4 points) Score 1. Annie is considering opening a certificate of deposit (CD) account. The CD account will earn interest. However, it does not allow her to deposit or withdraw money for 2 years. First Bank is offering a CD account with a 1.8% interest rate each month. The interest is compounded. In other words, the interest earned each month will be based on the total account balance from the previous month. Annie has $500 to start the account.

OpenStudy (ittybitty14):

(a) Fill in the table showing the amount of money in the CD account over 4 months. Banks always round down to the nearest cent when necessary. (b) If Annie graphed the function, what would the y-intercept be? How do you know this? (c) Is the relation a function? Explain why or why not. (d) Is the relation linear? If so, state the rate of change and write the equation of the relation, where x is the number of months and y is the amount in the account. If not, explain why not. Answer: Months 0 1 2 3 4 Amount in CD account $500.00

OpenStudy (will.h):

Normally for this kind of questions we use Exponential functions to calculate the interests rate or the amout that you will recieve after specific time for a certain interest rate. The Exponential function rule is: \[f(x)= p(1+r)^x\] where p= the inital amount of money and it is also the y intercept And R= The interest rate X= the amount of time. Therefore in your question the p= 500 R= 1.8% which means 1.8/100 which equals 0.018 And x=2 Now plug all these numbers in the function above so that we can find how much will she/he get after 2 years. \[f(x) = 500(1+ 0.018) ^2\] \[f(x) = 518.162\] That's the amount she will get after 2 years Now give me a moment for the questions A,B,C

OpenStudy (will.h):

are you good with the above info or do you need to ask about something

OpenStudy (ittybitty14):

Yeah I understand what you are saying thanks so much for helping me out!

OpenStudy (will.h):

No problem, am working on the rest,

OpenStudy (ittybitty14):

Okay! :))

OpenStudy (will.h):

okay so Part A we will have to work on linear functions, is that what your Assignment about or do you have to only work with Exponential functions?